OpenStudy (anonymous):

Write each statement as an equation, and find the number. If Twenty-Four is subtracted from a number, the result is the same as the number divided by nine. I Need to show all the Steps/work. I Am Horrible At Math Can Someone Please Take Me Through The Steps to get positive 27 For The Answer

OpenStudy (anonymous):

tryyyyyyyyyyyyyyyyy

OpenStudy (anonymous):

this is what i have so far \[x-24=\frac{ x }{ 9 }\]

OpenStudy (anonymous):

you would add the 24 to both sides correct

OpenStudy (johnweldon1993):

well...try multipling both sides by 9

OpenStudy (anonymous):

x-24(9)=\[\frac{ x }{ 9 }(9)\]

OpenStudy (anonymous):

correct?

OpenStudy (johnweldon1993):

right.....so 9(x-24) becomes....?

OpenStudy (anonymous):

9x-216

OpenStudy (johnweldon1993):

correct....and x/9 times 9 becomes?

OpenStudy (anonymous):

9x-219=\[\frac{ x }{ 81}\]

OpenStudy (anonymous):

correct?

OpenStudy (johnweldon1993):

no..... \[\frac{ x }{ 9 } * 9\] remember 9...is the same as 9/1 right? so you multiply straigh across x times 9...over 9 times 1 is 9x / 9 which simplifys to...?

OpenStudy (anonymous):

9x

OpenStudy (anonymous):

no 81 x

OpenStudy (johnweldon1993):

9 divided by 9....is 1.....so 1x....or just x...right?

OpenStudy (anonymous):

9x-216=x

OpenStudy (johnweldon1993):

correct now...?

OpenStudy (anonymous):

subtract 9x to both sides

OpenStudy (johnweldon1993):

well yeah you could do that...so you get.....?

OpenStudy (anonymous):

-216=-8x

OpenStudy (johnweldon1993):

and finally?

OpenStudy (anonymous):

subtract 8x to both sides which gives you x+27

OpenStudy (anonymous):

Thanks For the help i will give you a medal for your help

OpenStudy (johnweldon1993):

nooooo.....lol.....when you're down to -216 = -8x you divide both sides by -8.....you want x completely by itself

OpenStudy (anonymous):

i did that sorry i typed it wrong on here

OpenStudy (johnweldon1993):

ohh okay good! lol then good job :)